OpenStudy (user917):

Anyone here who is studying Real Analysis, can help me in a sequence exercise? Thank you

OpenStudy (zzr0ck3r):

sup

OpenStudy (abb0t):

What is your question?

OpenStudy (user917):

Hey @abb0t @zzr0ck3r , what is the upper and lower bound of these finite sum : {1/radical n} , {1/n^3} and {1/(1+n^3)}

OpenStudy (user917):

I think they all have 1 as an upper bound and 0 as a lower bound, is that correct?

OpenStudy (zzr0ck3r):

\[\frac{1}{\sqrt{n}}\]?

OpenStudy (user917):

yes

OpenStudy (zzr0ck3r):

then yes the upper bound would be 1, this happens when n=1. and as we let n grow the ratio gets smaller because n>1 then sqrt(n) > 1 for n natural.

OpenStudy (zzr0ck3r):

im with you on \[\frac{1}{n^3}\]is agin 1 when n = 1 and will get smaller as n grows

OpenStudy (zzr0ck3r):

but \[\frac{1}{1+n^3}\]the biggest this will get is 1/2

OpenStudy (zzr0ck3r):

\[2\le1+n^3\]for all n \[\implies\frac{1}{1+n^3}\le\frac{1}{2}\]

OpenStudy (zzr0ck3r):

and yes 0 for all the lower bounds

OpenStudy (user917):

wow thank you very much, is there any method which I should use to proof or I just need to say this and its fine? Thank you again

OpenStudy (zzr0ck3r):

did it say "prove the upper bound" or "find the upper bound'

OpenStudy (zzr0ck3r):

also , there are infinite upper bounds but 1 is the supremum (the smallest upper bound).

OpenStudy (zzr0ck3r):

it really depends on the teacher as far as proving these for the first one \[0\le1\le n^2\text{ for all n in N}\\0\le\frac{1}{n^2}\le1\text{ because }n>0\]

OpenStudy (zzr0ck3r):

this shows that an upper bound is 1 and a lower bound is 0 but if you need to show that they are the supremum and infimum then you need to do a little more.

OpenStudy (user917):

ok thank you :D

OpenStudy (user917):

@zzr0ck3r oh wait we need the lower and upper bound for the finite sum 1+1 + 1+ √2+1/√3 +···+ 1/√n

OpenStudy (user917):

so a lower bound is 1 and an upper bound is n

OpenStudy (zzr0ck3r):

what does this have to do with the last one?

OpenStudy (zzr0ck3r):

i dont see 1/n^3 or 1/(1+n^3)

OpenStudy (user917):

its not a sequence its a finite sum

OpenStudy (user917):

for example 1/1 + 1/2^3 + .... + 1/n^3

OpenStudy (zzr0ck3r):

ahh you have 3 different finite sums

OpenStudy (zzr0ck3r):

well then n^2 would be an upper bound because each term is amller than n, and we have n terms, so n*n=n^2

OpenStudy (zzr0ck3r):

n might work al well, but there are infinite answers so....

OpenStudy (user917):

thank you